OpenStudy (anonymous):

x^2-x-6/x^2-2x-3 divide by x^2+x-12/x^2+5x+4 Divide and simplify !! Show work please

OpenStudy (anonymous):

(x^2-x-6)*(x^2+5x+4)/(x^2+x-12)*(x^2-2x-3) = (x^4+5x^3+4x^2+5x^3-5x^2-4x-6x^2-30x-24)/(x^4-2x^3-3x^2+x^3-2x^2-3x-12x^2+24x+36) = (x^4+10x^3+3x^2-34x-24)/(x^4-x^3-17x^2+21x+36)